# A Call for New Recipes to Enhance Spatial Reasoning in MLLMs

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive performance in general vision-language tasks. However, recent studies have exposed critical limitations in their spatial reasoning capabilities. This deficiency in spatial reasoning significantly constrains MLLMs’ ability to interact effectively with the physical world, thereby limiting their broader applications. We argue that spatial reasoning capabilities will not naturally emerge from merely scaling existing architectures and training methodologies. Instead, this challenge demands dedicated attention to fundamental modifications in the current MLLM development approach. In this position paper, we first establish a comprehensive framework for spatial reasoning within the context of MLLMs. We then elaborate on its pivotal role in real-world applications. Through systematic analysis, we examine how individual components of the current methodology—from training data to reasoning mechanisms—influence spatial reasoning capabilities. This examination reveals critical limitations while simultaneously identifying promising avenues for advancement. Our work aims to direct the AI research community’s attention toward these crucial yet underexplored aspects. By highlighting these challenges and opportunities, we seek to catalyze progress toward achieving human-like spatial reasoning capabilities in MLLMs.
Multimodal Large Language Model, Spatial Reasoning
[…]
tcb@breakable
@@ 1 Introduction
Recent advancement in Large Language Models (LLMs) (Achiam et al., 2023; Dubey et al., 2024) and Multimodal Large Language Models (MLLMs) (Alayrac et al., 2022; Chen et al., 2024b; Wang et al., 2024b) has heightened discussions regarding artificial systems achieving human-level intelligence (Gignac & Szodorai, 2024). While extensive research has evaluated these models’ capabilities in specific domains, such as code generation (Jimenez et al., 2024), mathematical reasoning (Mirzadeh et al., 2024), and other tasks (Fu et al., 2023; Zhang et al., 2024a), demonstrating remarkable proficiency in specific areas, a fundamental component of general intelligence has been largely overlooked (Bubeck et al., 2023): spatial reasoning.
@@
In cognitive science (Waller & Nadel, 2013) and psychometrics (McGee, 1979), spatial reasoning (or namely spatial thinking) measures the capability to recall, generate, manipulate, and reason about spatial relations (Gilligan-Lee et al., 2022). Within Salthouse’s hierarchical structure (Salthouse, 2004), spatial reasoning emerges as an essential component alongside general reasoning capabilities in determining overall intelligence (Deary et al., 2010).
@@
However, within the machine learning and deep learning community, systematic investigation of language models’ spatial reasoning capabilities remains limited. While some studies have examined models’ ability to process textual spatial relations (Shi et al., 2022; Mirzaee & Kordjamshidi, 2022; Li et al., 2024b), their scope is constrained compared to the comprehensive framework established in cognitive science. Recent work with MLLMs has begun exploring this domain more extensively (Wang et al., 2024a; Zhang et al., 2024b; Ray et al., 2024). However, the field lacks a precise definition of spatial reasoning capabilities for these models, impeding rigorous analysis of model behavior and systematic improvement efforts.
@@
To this end, we deliver this position paper to discuss and provide viable pathways in how we can enhance spatial reasoning for MLLMs and hopefully inspire the advancement and traction in this area. The emergence of increasingly powerful MLLMs makes this investigation particularly timely, as these models can now process and integrate multimodal information—a prerequisite for human-like spatial reasoning. Enhanced spatial reasoning capabilities also contribute to unlocking the potential of MLLMs in downstream tasks and real-world applications that necessitate spatial awareness. Viewing from a higher level, the impact of advancing spatial reasoning in MLLMs extend far beyond incremental technological improvements. It transits from models that passively consume information to systems that can dynamically understand, predict, and interact with physical environments in ways that approach human-like cognitive flexibility. However, current MLLMs—from open-source to close-source models—consistently struggle with spatial reasoning tasks that are easy and straightforward for human (Ray et al., 2024; Zhang et al., 2024b; Yang et al., 2024). This introduces our discussion in this paper: contrary to the assumption that spatial reasoning emerges naturally as a fundamental cognitive skill, we assert that dedicated efforts are necessary to address this challenge.
@@
In this position paper, we argue that spatial reasoning demands greater attention from the research community and requires new recipes for enhancement in MLLMs. In Section 2, we first define two basic MLLM capabilities for spatial reasoning: relational reasoning and transformation reasoning, which extend beyond general visual understanding. We systematically analyse and identify several key factors that may influence the performance and advancement of MLLMs in spatial reasoning through theoretical analysis and evidence from previous work. As illustrated in Figure 1, the factors include but are not limited to the data and objectives during training (Section 3.1 and 3.3), model architectures, reasoning strategies during inference (Section 3.2 and 5), etc. These components lack the cultivation and design of MLLMs tailored for spatial reasoning, thereby limiting their performance. We support our arguments with a detailed case study in Section 6 and explore alternative perspectives in Section 7 to stimulate further discussion.
@@ 2 Spatial Reasoning
Spatial Reasoning is fundamental to human cognition, encompassing the ability to understand and reason about the spatial relationships among objects, their movements, and interactions with the environment (Chabris et al., 2006; Newcombe, 2024). This sophisticated cognitive skill relies on complex neural mechanisms that integrate visual perception, working memory, executive function, and spatial transformation abilities (Waller & Nadel, 2013; Knauff et al., 2002). While humans naturally develop this capability, achieving this functionality in AI systems presents unique challenges that demand our immediate attention.
@@
In this paper, we focus on spatial reasoning informed by visual cues, a crucial capability associated with visual-spatial intelligence (Yang et al., 2024). Drawing upon established frameworks from cognitive psychology (Yılmaz, 2009) and recent advances in MLLMs (Yang et al., 2024), we propose a comprehensive framework for spatial reasoning in the context of MLLMs. Specifically, we decompose spatial reasoning into two fundamental capabilities:
@@
Relational Reasoning encompasses the ability to understand and identify spatial relationships between objects, such as relative positions, orientations, and perspective (Chen et al., 2024a; Cheng et al., 2024; Zhang et al., 2024b). This capability captures the static aspect of spatial reasoning, focusing on analyzing spatial configurations in a given context.
@@
Transformation Reasoning involves the ability to reason about spatial transformations, including rotations, translations, perspective changes, and other dynamic manipulations of spatial configurations (Ray et al., 2024; Zhang et al., 2024b; Yang et al., 2024). This capability addresses the dynamic aspect of spatial reasoning.
@@
By defining these two capabilities, we aim to cover the static and dynamic dimensions of spatial reasoning comprehensively. Relational reasoning captures the foundational understanding of “what is where,” while transformation reasoning extends this to “how things change or could change” within spatial contexts. These dimensions closely correspond to the well-established spatial abilities in cognitive psychology: spatial relations and spatial visualization (Yılmaz, 2009). This alignment not only adapts cognitive theory to the context of MLLMs but also demonstrates its completeness in capturing the essential components of spatial reasoning.
@@
These spatial reasoning capabilities encompass both 2D (Liu et al., 2023) and 3D scenes (Hong et al., 2023), and are fundamental to a wide spectrum of real-world applications where MLLMs interface with spatial information. In Embodied AI, Relational Reasoning enables robots to understand spatial relationships between objects, such as recognizing obstacles and determining object positions. While simultaneously employing Transformation Reasoning, robots can navigate, manipulate objects, and interpret spatial instructions (Hao et al., 2024; Ishida, 2024). For Autonomous Driving, spatial reasoning capabilities enable MLLMs to perceive and respond to dynamic environments (Guo et al., 2024). In Remote Sensing applications, spatial reasoning facilitates the analysis of satellite imagery for urban planning, environmental monitoring, and geographical feature extraction (Zhan et al., 2024; Li et al., 2024c). Within Medical Imaging, both Relational and Transformation Reasoning capabilities are crucial for tasks such as anatomical structure localization and surgical planning from 3D medical scans (Xia et al., 2024b, a). In educational applications, MLLMs equipped with spatial reasoning capabilities can assist learners in understanding spatial concepts by leveraging Relational Reasoning for geometric relationships and Transformation Reasoning for dynamic manipulations like rotations (Xing et al., 2024; Ramakrishnan et al., 2024). These diverse applications underscore the critical role of spatial reasoning in enabling MLLMs to understand and interact with the physical world in meaningful ways.
@@
Building upon our framework of spatial reasoning, we conduct a systematic analysis of the current components of MLLM, identify their specific limitations and challenges in spatial reasoning, and propose opportunities for improvement in the following sections. Our goal is to catalyze research efforts that will advance MLLMs’ spatial reasoning abilities to meet the demands of real-world applications.
@@ 3.1 Training Data
Pre-training Data. Pre-training data mainly has two critical functions: aligning different modalities and providing world knowledge (Yin et al., 2023). Recent analyses (Kamath et al., 2023) have revealed limitations in pre-training datasets on spatial relationships, particularly the insufficient spatial annotations. Building on the research of Kamath et al. examining spatial prepositions in LAION-2B (Schuhmann et al., 2022), we expanded this investigation to examine additional prominent datasets, include CC-3M (Sharma et al., 2018), SBU Captions (Ordonez et al., 2011), and ShareGPT4V-PT (Chen et al., 2025). We carefully filtered out spatial prepositions used in non-spatial contexts, such as “under \$25.”
@@
As shown in Tab 1, our findings indicate that LAION-2B, CC-3M, and SBU Captions contain notably low proportions of spatial prepositions in their captions. This indicates a great gap in their coverage of basic spatial relationships, which significantly constrains the Relational Reasoning capability of MLLMs. While ShareGPT4V-PT offers more detailed spatial descriptions, its high-quality annotations come at substantial cost. Furthermore, the limitations become increasingly evident in the context of Transformation Reasoning capability, as current datasets rarely contain descriptions of spatial transformations. For instance, in What’s UP (Kamath et al., 2023), the performance was improved by leveraging specialized spatial relation corpora based on existing datasets, such as COCO (Lin et al., 2014). However, fundamental challenges such as perspective transformation remain unresolved by current datasets.
@@
Beyond basic spatial relation descriptions, other datasets, such as XVLM (Zeng et al., 2022), have incorporated denser supervision signals like bounding box. But existing approaches remain constrained by their reliance on natural language for annotating spatial information in images. Text descriptions primarily capture unary (single object) or binary (two object) relationships (Comsa & Narayanan, 2023), such as “the cup is on the table” or “the person stands between two chairs”. This fundamental limitation makes it inefficient to represent complex spatial configurations involving multiple objects and their intricate spatial relationships. When describing sophisticated layouts, natural language descriptions become verbose and potentially ambiguous (Landau & Lakusta, 2009). Looking ahead, we believe the field should explore more sophisticated spatial representation schemes that transcend these limitations. Alternative approaches like topological representations could offer more compact and precise ways to encode spatial relationships, potentially revolutionizing how we represent spatial information in MLLMs.
Fine-tuning Data for Spatial Reasoning. Recent work has made significant progress in developing fine-tuning datasets that target different aspects of spatial reasoning capabilities. We analyze these advances through the lens of both Relational and Transformation Reasoning abilities.
@@
SpatialVLM (Chen et al., 2024a) performs depth estimation to lift internet-scale images to object-centric 3D point clouds and use it to synthesize VQA data embedded with 3D spatial reasoning supervision. SpatialRGPT (Cheng et al., 2024) uses 3D scene graphs to construct template-based spatial descriptions. These descriptions, combined with region tags are utilized as input. Complementary to these datasets focusing on quantitative relationships like distances or size difference, the recent work SAT (Ray et al., 2024) extended the scope to dynamic spatial transformation data. Through targeting the reasoning of spatial transformations and other complex tasks, SAT not only improved performance on dynamic spatial tasks but also enhanced zero-shot capabilities on existing real-image spatial benchmarks.
@@ 3.2 Model Architecture
Vision Encoder. The vision encoder serves as the primary interface for extracting spatial information from visual inputs. Current vision encoders, such as Vision Transformers (ViTs) (Dosovitskiy et al., 2021), predominantly adopt transformer-based architectures. Research in experimental psychology and neuroscience has shown that place and grid cells in the hippocampus and entorhinal cortex contribute to spatial encoding, helping organisms perceive and remember locations (Moser et al., 2008). Drawing parallels to artificial neural systems, positional embeddings in vision encoders serve a similar fundamental purpose. Since the self-attention mechanism in transformer is independent of the token index or positions, the vision encoder depends on position embeddings to provide essential spatial information.
@@
There are two primary methods in position embedding for Vision Transformer: Absolute Positional Embedding (APE) (Dosovitskiy et al., 2021; Devlin et al., 2019) and Relative Position Embedding (RPE) (Liu et al., 2021; Shaw et al., 2018). APE encodes the absolute position of each token through sinusoidal functions or learnable embeddings, reminiscent of grid cells. RPE, similar to place cells, integrates relative position biases into the attention matrix, modeling the relative spatial relationships between tokens.
@@
Nevertheless, both APE and RPE present certain limitations. APE lacks flexibility and the ability for extrapolation, whereas RPE restricts the attention module, making it challenging to reflect sufficient positional information to the output (Luo et al., 2022). Therefore, in the visual encoder, APE may struggle to handle spatial transformations such as rotation and translation. Similarly, RPE could limits the visual encoder’s ability to fully encode and propagate complex spatial information, particularly when dealing with precise distance estimation. Swin Transformer (Liu et al., 2021) explored a simple approach to combine RPE and APE, but results were not promising. Inspired by the coordinated function of grid and place cells in human cognition, it may be worthwhile to investigate more advanced hybrid positional encoding methods that could better capture and integrate spatial information for improved performance.
@@
Connector. The connector, when present, plays a crucial role in preserving spatial information while bridging different modalities. It must maintain spatial fidelity while projecting visual features into the language space. Specifically, the visual features $f\in\mathbb{R}^{P\times d_{v}}$ from encoders are processed by the connector $C$, where $P$ is the number of visual patches and $d_{v}$ is the channel dimension. The alignment can be formulated as $x=C(f),C:\mathbb{R}^{P\times d_{v}}\rightarrow\mathbb{R}^{Q\times D}$. Here, $x\in\mathbb{R}^{Q\times D}$ represents the visual tokens that are input into the LLM, where,$Q$ is the numbder of visual tokens and $D$ is the hidden size of LLM. In a recent study on connector (Lin et al., 2024), the authors categorize connectors into two types: feature-preserving connectors, where $P=Q$, and feature-compressing connectors, where $P>Q$.
@@
The feature-preserving connectors can retain detailed information in the visual tokens (Lin et al., 2024). Llava series (Liu et al., 2024) are one of the popular works in this category. While feature-preserving connectors focus on retaining detailed visual information, feature-compressing connectors reduce feature patches for better computational efficiency without losing essential information. Various approaches have been proposed to achieve this balance: Emu2 (Sun et al., 2024a) employs local average pooling, and BLIP-2 (Li et al., 2023) utilizes the Q-Former, a cross-attention connector with fixed learnable queries for global weighted pooling. Besides, HoneyBee (Cha et al., 2024) introduces the C-Abstractor, leveraging convolutional networks for local weighted pooling. However, the complex connectors like Q-Former may require more aligned data for better results (Lin et al., 2024). Furthermore,the researchers from NVLM (Dai et al., 2024) have found that the cross-attention to latent array in the Perceiver (Jaegle et al., 2021; Alayrac et al., 2022) mixes the input image tokens, potentially disrupting the spatial relationships between image patches, which are crucial for spatial reasoning.
@@
These connector designs face a fundamental trade-off between preserving spatial information and computational efficiency. Recent work has started to address this challenge - notably, Cambrian (Tong et al., 2024) proposed a novel Spatial Vision Aggregator design that maintains spatial structure during cross-attention operations. This innovation demonstrates the potential for architectural improvements in this critical component. We advocate for future research to develop novel connector architectures that can better balance this trade-off while explicitly preserving spatial information.
@@
The prevalent approach in current MLLMs, which flattens visual tokens into a 1D sequence before applying RoPE-1D (Su et al., 2024) alongside text tokens in the LLM backbone, inherently introduces influence to spatial information. This linearization of 2D visual information fundamentally compromises the spatial structure of images. Recent work like Qwen2-VL (Wang et al., 2024b) has made promising progress in addressing this limitation through M-RoPE, which decomposes positional encoding into temporal, height, and width components. The superior performance brought by M-RoPE on various multimodal tasks suggests that proper position embedding design can significantly impact the model’s ability, including spatial reasoning. Therefore, future work should explore better position embeddings along this path to enhance spatial reasoning in MLLMs.
In the backbone, various attention masks can be utilized to control the information flow from different modalities. A typical one is the causal mask, which requires each token to only depend on prior context for generation. However, this approach may constrain the model’s ability to process spatial relationships holistically, as visual spatial information inherently requires bidirectional attention patterns. Recent works (Beyer et al., 2024; Steiner et al., 2024) have demonstrated improved performance on spatial reasoning benchmark VSR (Liu et al., 2023) by allowing more tokens to participate in the reasoning process. This suggests that the uniform application of causal masking may be suboptimal. Therefore, we encourage future studies to explore attention masking strategies that better align with the non-sequential nature of spatial information in visual tokens.

### v2

# Scaling and Beyond: Advancing Spatial Reasoning in MLLMs Requires New Recipes

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive performance in general vision-language tasks. However, recent studies have exposed critical limitations in their spatial reasoning capabilities. This deficiency in spatial reasoning significantly constrains MLLMs’ ability to interact effectively with the physical world, thereby limiting their broader applications. We argue that spatial reasoning capabilities will not naturally emerge from merely scaling existing architectures and training methodologies. Instead, this challenge demands dedicated attention to fundamental modifications in the current MLLM development approach. In this position paper, we first establish a comprehensive framework for spatial reasoning within the context of MLLMs. We then elaborate on its pivotal role in real-world applications. Through systematic analysis, we examine how individual components of the current methodology, from training data to reasoning mechanisms, influence spatial reasoning capabilities. This examination reveals critical limitations while simultaneously identifying promising avenues for advancement. Our work aims to direct the AI research community’s attention toward these crucial yet underexplored aspects. By highlighting these challenges and opportunities, we seek to catalyze progress toward achieving human-like spatial reasoning capabilities in MLLMs.
@@ 1 Introduction
Recent advancement in Large Language Models (LLMs) [1, 2] and Multimodal Large Language Models (MLLMs) [3, 4, 5] has heightened discussions regarding artificial systems achieving human-level intelligence [6]. While extensive research has evaluated these models’ capabilities in specific domains, such as code generation [7], mathematical reasoning [8], and other tasks [9, 10], demonstrating remarkable proficiency in specific areas, a fundamental component of general intelligence, spatial reasoning, remains substantially underdeveloped [11, 12, 13].
@@
In cognitive science [14] and psychometrics [15], spatial reasoning (or namely spatial thinking) measures the capability to recall, generate, manipulate, and reason about spatial relations [16]. Within Salthouse’s hierarchical structure [17], spatial reasoning emerges as an essential component alongside general reasoning capabilities in determining overall intelligence [18].
@@
However, within the machine learning and deep learning community, systematic investigation of language models’ spatial reasoning capabilities remains limited. While some studies have examined models’ ability to process textual spatial relations [19, 20, 21], their scope is constrained compared to the comprehensive framework established in cognitive science. Recent work with MLLMs has begun exploring this domain more extensively [22, 23, 11, 24]. Nevertheless, the field lacks a precise definition of spatial reasoning capabilities for these models, impeding rigorous analysis of model behavior and systematic improvement efforts.
@@
To this end, we deliver this position paper to discuss and provide viable pathways in how we can enhance spatial reasoning for MLLMs and hopefully inspire the advancement and traction in this area. The emergence of increasingly powerful MLLMs makes this investigation particularly timely, as these models can now process and integrate multimodal information—a prerequisite for human-like spatial reasoning. Enhanced spatial reasoning capabilities also contribute to unlocking the potential of MLLMs in downstream tasks and real-world applications that necessitate spatial awareness. Viewing from a higher level, the impact of advancing spatial reasoning in MLLMs extend far beyond incremental technological improvements. It transits from models that passively consume information to systems that can dynamically understand, predict, and interact with physical environments in ways that approach human-like cognitive flexibility. However, current MLLMs—from open-source to closed-source models—consistently struggle with spatial reasoning tasks that are easy and straightforward for human [11, 23, 13]. This introduces our discussion in this paper: contrary to the assumption that spatial reasoning emerges naturally as a fundamental cognitive skill, we assert that dedicated efforts are necessary to address this challenge.
@@
In this position paper, we argue that advancing spatial reasoning in MLLMs requires more than scaling; it requires new design principles and targeted innovations. In Section 2, we first define two basic MLLM capabilities for spatial reasoning: relational reasoning and transformation reasoning, which extend beyond general visual understanding. We systematically analyze and identify several key factors that may influence the performance and advancement of MLLMs in spatial reasoning through theoretical analysis and evidence from previous work. As illustrated in Figure 1, the factors include but are not limited to the data and objectives during training (Section 3.1 and 3.3), model architectures, reasoning strategies during inference (Section 3.2 and 5), etc. These components lack the cultivation and design of MLLMs tailored for spatial reasoning, thereby limiting their performance. Addutionally, We support our arguments with a detailed case study in Section 6.
@@ 2 Spatial Reasoning
Spatial Reasoning is fundamental to human cognition, encompassing the ability to understand and reason about the spatial relationships among objects, their movements, and interactions with the environment [25, 26]. This sophisticated cognitive skill relies on complex neural mechanisms that integrate visual perception, working memory, executive function, and spatial transformation abilities [14, 27]. While humans naturally develop this capability, achieving this functionality in AI systems presents unique challenges that demand our immediate attention.
@@
In this paper, we focus on spatial reasoning informed by visual cues, a crucial capability associated with visual-spatial intelligence [13]. Drawing upon established frameworks from cognitive psychology [28] and recent advances in MLLMs [13], we propose a comprehensive framework for spatial reasoning in MLLMs. Specifically, we decompose it into two fundamental capabilities:
@@
Relational Reasoning encompasses the ability to understand and identify spatial relationships between objects, such as relative positions, orientations, and perspective [29, 30, 23]. This capability captures the static aspect of spatial reasoning, focusing on analyzing spatial configurations.
@@
Transformation Reasoning involves the ability to reason about spatial transformations, including rotations, translations, perspective changes, and other dynamic manipulations of spatial configurations [11, 23, 13]. This capability addresses the dynamic aspect of spatial reasoning.
@@
By defining these two capabilities, we aim to cover the static and dynamic dimensions of spatial reasoning comprehensively. Relational reasoning captures the foundational understanding of “what is where," while transformation reasoning extends this to “how things change or could change" within spatial contexts. These dimensions closely correspond to the well-established spatial abilities in cognitive psychology: spatial relations and spatial visualization [28]. This alignment not only adapts cognitive theory to the context of MLLMs but also demonstrates its completeness in capturing the essential components of spatial reasoning.
@@
These spatial reasoning capabilities encompass both 2D [31] and 3D scenes [32], and are fundamental to a wide spectrum of real-world applications where MLLMs interface with spatial information. In Embodied AI, Relational Reasoning enables robots to understand spatial relationships between objects, such as recognizing obstacles and determining object positions. While simultaneously employing Transformation Reasoning, robots can navigate, manipulate objects, and interpret spatial instructions [33, 34]. For Autonomous Driving, spatial reasoning capabilities enable MLLMs to perceive and respond to dynamic environments [35]. In Remote Sensing applications, spatial reasoning facilitates the analysis of satellite imagery for urban planning, environmental monitoring, and geographical feature extraction [36, 37]. Within Medical Imaging, both Relational and Transformation Reasoning capabilities are crucial for tasks such as anatomical structure localization and surgical planning from 3D medical scans [38, 39]. In educational applications, MLLMs equipped with spatial reasoning capabilities can assist learners in understanding spatial concepts by leveraging Relational Reasoning for geometric relationships and Transformation Reasoning for dynamic manipulations like rotations [40, 41]. These diverse applications underscore the critical role of spatial reasoning in enabling MLLMs to understand and interact with the physical world in meaningful ways.
@@
Building upon our framework of spatial reasoning, we conduct a systematic analysis of the current components of MLLM, identify their specific limitations and challenges in spatial reasoning, and propose opportunities for improvement in the following sections. Our goal is to catalyze research efforts that will advance MLLMs’ spatial reasoning abilities to help real-world applications.
@@ 3.1 Training Data
Pre-training Data. Pre-training data mainly has two critical functions: aligning different modalities and providing world knowledge [42]. Recent analyses [43] have revealed limitations in pre-training datasets on spatial relationships, particularly the insufficient spatial annotations. Building on What’s UP, we extend the analysis to additional prominent datasets, including LAION-2B [44], CC-3M [45], SBU Captions [46], and ShareGPT4V-PT [47]. We carefully filtered out spatial prepositions used in non-spatial contexts, such as "under \$25."
@@
As shown in Tab 1, our findings indicate that LAION-2B, CC-3M, and SBU Captions contain notably low proportions of spatial prepositions in their captions. This indicates a great gap in their coverage of basic spatial relationships, which significantly constrains the Relational Reasoning capability of MLLMs. While ShareGPT4V-PT offers more detailed spatial descriptions, its high-quality annotations come at substantial cost. Furthermore, the limitations become increasingly evident in the context of Transformation Reasoning capability, as current datasets rarely contain descriptions of spatial transformations. For instance, in What’s UP [43], the performance was improved by leveraging specialized spatial relation corpora based on existing datasets, such as COCO [48]. However, fundamental challenges such as perspective transformation remain unresolved.
@@
Beyond basic spatial relation descriptions, other datasets, such as XVLM [49], have incorporated denser supervision signals like bounding box. But existing approaches remain constrained by their reliance on natural language for annotating spatial information in images. Text descriptions primarily capture unary (single object) or binary (two object) relationships [50], such as “the cup is on the table" or “the person stands between two chairs". This fundamental limitation makes it inefficient to represent complex spatial configurations involving multiple objects and their intricate spatial relationships. When describing sophisticated layouts, natural language descriptions become verbose and potentially ambiguous [51]. Looking ahead, we believe the field should explore more sophisticated spatial representation schemes that transcend these limitations. Alternative approaches like topological representations could offer more compact and precise ways to encode spatial relationships, potentially revolutionizing how we represent spatial information in MLLMs.
@@
Fine-tuning Data for Spatial Reasoning. Recent work has made significant progress in developing fine-tuning datasets that target different aspects of spatial reasoning capabilities. SpatialVLM [29] performs depth estimation to lift internet-scale images to object-centric 3D point clouds and use it to synthesize VQA data embedded with 3D spatial reasoning supervision. SpatialRGPT [30] uses 3D scene graphs to construct template-based spatial descriptions. These descriptions, combined with region tags are utilized as input. While the above datasets focus on quantitative relationships such as distances and size differences, SAT [11] broadens the scope to dynamic spatial transformations, enhancing both task-specific performance and zero-shot generalization on real-image spatial benchmarks. Complementing these efforts, RoboSpatal [52] provides egocentric images and 3D scans with spatial annotations to enhance reasoning in robotic settings across diverse reference frames.
@@ 3.2 Model Architecture
Vision Encoder. The vision encoder serves as the primary interface for extracting spatial information from visual inputs. Current vision encoders, such as Vision Transformers (ViTs) [53], predominantly adopt transformer-based architectures. Research in experimental psychology and neuroscience has shown that place and grid cells in the hippocampus and entorhinal cortex contribute to spatial encoding, helping organisms perceive and remember locations [54]. Drawing parallels to artificial neural systems, positional embeddings in vision encoders serve a similar fundamental purpose. Since the self-attention mechanism in transformer is independent of the token index or positions, the vision encoder depends on position embeddings to provide essential spatial information.
@@
There are two primary methods in position embedding for Vision Transformer: Absolute Positional Embedding (APE) [53, 55] and Relative Position Embedding (RPE) [56, 57]. APE encodes the absolute position of each token through sinusoidal functions or learnable embeddings, reminiscent of grid cells. RPE, similar to place cells, integrates relative position biases into the attention matrix, modeling the relative spatial relationships between tokens.
@@
Nevertheless, both APE and RPE present certain limitations. APE lacks flexibility and the ability for extrapolation, whereas RPE restricts the attention module, making it challenging to reflect sufficient positional information to the output [58]. Therefore, in the visual encoder, APE may struggle to handle spatial transformations such as rotation and translation. Similarly, RPE could limit the visual encoder’s ability to fully encode and propagate complex spatial information, particularly when dealing with precise distance estimation. Swin Transformer [56] explored a simple approach to combine RPE and APE, but results were not promising. Inspired by the coordinated function of grid and place cells in human cognition, it may be worthwhile to investigate more advanced hybrid positional encoding methods that could better capture and integrate spatial information.
In parallel, recent studies have explored the use of pre-trained 3D encoders [59, 60, 61, 62]. However, these 3D encoders are trained on domain-specific tasks [63, 64, 65] and lack the general visual-language understanding ability like CLIP. This mismatch introduces a trade-off between generalization and 3D spatial understanding in MLLMs. A common strategy to mitigate this issue is to adopt a dual-branch encoder design [66], which integrates both a general-purpose visual encoder and a 3D encoder. Nonetheless, this approach often leads to feature redundancy and potential representation entanglement, limiting its scalability and efficiency.
@@
Connector. The connector, when present, plays a crucial role in preserving spatial information while bridging different modalities. It must maintain spatial fidelity while projecting visual features into the language space. Specifically, the visual features $f\in\mathbb{R}^{P\times d_{v}}$ from encoders are processed by the connector $C$, where $P$ is the number of visual patches and $d_{v}$ is the channel dimension. The alignment can be formulated as $x=C(f),C:\mathbb{R}^{P\times d_{v}}\rightarrow\mathbb{R}^{Q\times D}$. Here, $x\in\mathbb{R}^{Q\times D}$ represents the visual tokens that are input into the LLM, $Q$ is the numbder of visual tokens and $D$ is the hidden size of LLM. In a recent study on connector [67], the authors categorize connectors into two types: feature-preserving connectors, where $P=Q$, and feature-compressing connectors, where $P>Q$.
@@
The feature-preserving connectors can retain detailed information in the visual tokens [67]. Llava series [68] are one of the popular works in this category. While feature-preserving connectors focus on retaining detailed visual information, feature-compressing connectors reduce feature patches for better computational efficiency without losing essential information. Various approaches have been proposed to achieve this balance: Emu2 [69] employs local average pooling, and BLIP-2 [70] utilizes the Q-Former, a cross-attention connector with fixed learnable queries for global weighted pooling. Besides, HoneyBee [71] introduces the C-Abstractor, leveraging convolutional networks for local weighted pooling. However, the complex connectors like Q-Former may require more aligned data for better results [67]. Furthermore, the researchers from NVLM [72] have found that the cross-attention to latent array in the Perceiver [73, 3] mixes the input image tokens, potentially disrupting the spatial relationships between image patches, which are crucial for spatial reasoning.
@@
These connector designs face a fundamental trade-off between preserving spatial information and computational efficiency. Recent work has started to address this challenge - notably, Cambrian [74] proposed a novel Spatial Vision Aggregator design that maintains spatial structure during cross-attention operations. This innovation demonstrates the potential for architectural improvements in this critical component. We advocate for future research to develop novel connector architectures that can better balance this trade-off while explicitly preserving spatial information.
@@
The prevalent approach in current MLLMs, which flattens visual tokens into a 1D sequence before applying RoPE-1D [75] alongside text tokens in the LLM backbone, inherently introduces influence to spatial information. This linearization of 2D visual information fundamentally compromises the spatial structure of images. Recent work like Qwen2-VL [5] has made promising progress in addressing this limitation through M-RoPE, which decomposes positional encoding into temporal, height, and width components. The superior performance brought by M-RoPE on various multimodal tasks suggests that proper position embedding design can significantly impact the model’s ability, including spatial reasoning. Therefore, future work should explore better position embeddings along this path to enhance spatial reasoning in MLLMs.
In the backbone, various attention masks can be utilized to control the information flow from different modalities. A typical one is the causal mask, which requires each token to only depend on prior context for generation. However, this approach may constrain the model’s ability to process spatial relationships, as visual spatial information inherently requires bidirectional attention patterns. Recent works [76, 77] have demonstrated improved performance on spatial reasoning benchmark VSR [31] by allowing more tokens to participate in the reasoning process. This suggests that the application of causal masking may be suboptimal. Therefore, we encourage future studies to explore attention masking strategies that better align with the non-sequential nature of visual spatial information.